# Approximate Nash Equilibria Algorithms for Network Design Games

### 05 Dec 2023 v1

@@ 3 Existence of $\alpha$-APNE
    $log_{2}(1+x+y)-log_{2}(1+x)\geq\frac{y}{x+y}$;
  2. 2.
    $log_{2}(1+x+y)-log_{2}(1+x)<log_{2}[e(1+y)]\cdot\frac{y}{x+y}$.
Proof
We first point out $(1+\frac{1}{x})^{x}$ increases monotonically and $\lim_{x\to\infty}(1+\frac{1}{x})^{x}=e$. For part 1, if $y\geq x+1$, then
$log_{2}(1+x+y)-log_{2}(1+x)=log_{2}(1+\frac{y}{x+1})\geq 1\geq\frac{y}{x+y}$
if $y<x+1$, then we have
$(1+\frac{y}{x+1})^{\frac{x+y}{y}}\geq(1+\frac{y}{x+1})^{\frac{x+1}{y}}\geq 2$
Therefore, we have $log_{2}(1+\frac{y}{x+1})\geq\frac{y}{x+y}$, which completes the proof for part 1.
For part 2, we have
$(1+\frac{y}{x+1})^{\frac{x+y}{y}}$ $=(1+\frac{y}{x+1})^{\frac{x+1}{y}}(1+\frac{y}{x+1})^{\frac{y-1}{y}}$
$<(1+\frac{y}{x+1})^{\frac{x+1}{y}}(1+\frac{y}{x+1})$
$\leq e(1+y)$
Similarly to the first part, $log_{2}(1+\frac{y}{x+1})\leq log_{2}[e(1+y)]\cdot\frac{y}{x+y}$, which completes the proof for part 2.
With the Lemma 1, we now prove Theorem 3.1.
Proof
For a weighted network design game, We define a function $\Phi$ as follows:
$\Phi(P)=\sum_{e\in E}c_{e}(W_{e})log_{2}(1+W_{e})$
Claim
$\Phi$ is an $\alpha$-approximate potential function for the weighted network design game.
To prove the claim, we need to show that $\epsilon$-improving deviations strictly decrease the function $\Phi$. Specifically, we consider an $\epsilon$-improving deviation of player $i$ from the profile $P=(P_{i},P^{\prime}_{-i})$ to the profile $P^{\prime}=(P^{\prime}_{i},P_{-i})$. Without loss of generality, we can assume that $P_{i}$ and $P^{\prime}_{i}$ are disjoint. Otherwise, we can replace $P_{i}$ and $P^{\prime}_{i}$ with $P_{i}\backslash P^{\prime}_{i}$ and $P^{\prime}_{i}\backslash P_{i}$ in the following discussion. By the definition of $\alpha$-improving, we have
$\alpha\cdot\sum_{e\in P^{\prime}_{i}}{c_{e}(W_{e})}\cdot\frac{w_{i}}{W_{e}+w_{i}}\leq\sum_{e\in P_{i}}{c_{e}(W_{e})}\cdot\frac{w_{i}}{W_{e}}.$
Here, $W_{e}$ denotes the total weight on edge $e$ before the deviation. Since $c_{e}(W_{e})=a_{e}W_{e}+b_{e}$, we have
$\alpha\cdot\sum_{e\in P^{\prime}_{i}}{(a_{e}(W_{e}+w_{i})+b_{e})}\cdot\frac{w_{i}}{W_{e}+w_{i}}\leq\sum_{e\in P_{i}}{(a_{e}W_{e}+b_{e})}\cdot\frac{w_{i}}{W_{e}}.$ (5)
Then we can derive the following:
$\Delta\Phi$ $=\sum_{e\in P^{\prime}_{i}}{(a_{e}W_{e}+b_{e})}(log_{2}(1+W_{e}+w_{i})-log_{2}(1+W_{e}))$
$\,\quad-\sum_{e\in P_{i}}{(a_{e}W_{e}+b_{e})}(log_{2}(1+W_{e})-log_{2}(1+W_{e}-w_{i}))$
$\,\quad+\sum_{e\in P^{\prime}_{i}}(a_{e}w_{i})log_{2}(1+W_{e}+w_{i})-\sum_{e\in P_{i}}(a_{e}w_{i})log_{2}(1+W_{e}-w_{i})$ (6)
$<\sum_{e\in P^{\prime}_{i}}{(a_{e}W_{e}+b_{e})}log_{2}[e(1+w_{i})]\frac{w_{i}}{W_{e}+w_{i}}$
$\,\quad-\sum_{e\in P_{i}}{(a_{e}W_{e}+b_{e})}\frac{w_{i}}{W_{e}}+\sum_{e\in P^{\prime}_{i}}(a_{e}w_{i})log_{2}(1+W_{e}+w_{i})$ (7)
$=\sum_{e\in P^{\prime}_{i}}(a_{e}w_{i})(log_{2}(1+W_{e}+w_{i})+log_{2}[e(1+w_{i})]\frac{w_{i}}{W_{e}+w_{i}})$
$\,\quad+\sum_{e\in P^{\prime}_{i}}b_{e}log_{2}[e(1+w_{i})]\frac{w_{i}}{W_{e}+w_{i}}-\sum_{e\in P_{i}}{(a_{e}W_{e}+b_{e})}\frac{w_{i}}{W_{e}}$ (8)
$<\alpha\sum_{e\in P^{\prime}_{i}}{(a_{e}(W_{e}+w_{i})+b_{e})}\cdot\frac{w_{i}}{W_{e}+w_{i}}-\sum_{e\in P_{i}}{(a_{e}W_{e}+b_{e})}\cdot\frac{w_{i}}{W_{e}}$ (9)
$=\alpha c_{i}(P^{\prime})-c_{i}(P),$ (10)
where $\alpha=2log_{2}(1+W+w_{max})$. Therefore, we have proved
$\Phi(P^{\prime})-\Phi(P)<\alpha c_{i}(P^{\prime})-c_{i}(P),$ (11)
which completes the proof of the claim. Moreover, since there are a finite number of profiles, the theorem can be directly derived from the claim.
4 Better Response Dynamics
Initially, we propose an algorithm to find the $\alpha$-APNE of the game $\Gamma$ as followings:
Algorithm 1 $\epsilon$-Approximate Better Response Dynamics($\epsilon$-ABRD)
Input:
A weighted network design game $\Gamma=(G,\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$; a strategy profile $P\in\mathcal{P}$
Output:
An $\alpha$-APNE of game $\Gamma$
  1. 1.
    While $P$ is not an $\alpha$-APNE do
@@ 4 Better Response Dynamics
Since computing $\alpha$-APNE is PLS-complete, we need to make further assumption on the game.
@@
Let $X$ be a random variable, we say $X$ is a $\phi$-smooth random variable if and only if $X$ is defined on $[0,1]$ and the cumulative distribution function of $X$ is upper bounded by a constant $\phi\geq 1$.
Before proving the algorithm above can fine the $\alpha$-APNE in a polynomial time, we first prove two lemmas as preparation.
@@
Let $X_{1},X_{2},...,X_{n};Y_{1},Y_{2},...,Y_{n}$ be $\phi$-smooth independent random variables on [0,1]. Let $g(x)=\min\{\frac{\alpha}{x}ln\frac{\alpha}{x},n^{\lambda}\}$. Then for any $\lambda\in N$ and $\alpha\geq 1$, we have
$\mathbb{E}[g(\min_{i\in\{1,2,...n\}}\{X_{i}+Y_{i}\})]\leq\alpha\phi n\cdot ln(\alpha\phi n)\cdot(\lambda+1)^{2}$
@@
For two random variables $S,T$, we define $S\preceq T$ if and only if $F_{T}(x)\leq F_{S}(x)$ for any $x\in\mathbb{R}$, where $F_{S}(x),F_{T}(x)$ are the cumulative distribution function of $S,T$ respectively. To prove the lemma, we define two sets of independent uniformly distributed random variables $S_{1},S_{2},...,S_{n};T_{1},T_{2},...,T_{n}$ on $[0,\frac{1}{\phi}]$.
@@
For any $i\in\{1,2,...,n\}$, we have
$min_{i\in\{1,2,...,n\}}S_{i}\preceq min_{i\in\{1,2,...,n\}}X_{i},min_{i\in\{1,2,...,n\}}T_{i}\preceq min_{i\in\{1,2,...,n\}}Y_{i}.$
@@
Since $S_{i}$ is defined on $[0,\frac{1}{\phi}]$, we have $F_{S_{i}}(t)=1$ for any $t\geq\frac{1}{\phi}$, then we have $F_{X_{i}}(t)\leq F_{S_{i}}(t)$ definitely. When $0\leq t\leq\frac{1}{\phi}$, since $X_{i}$ is a $\phi$-smooth random variable for any $i\in\{1,2,...,n\}$, its probability density function is upper-bounded by $\phi$. Then we have
@@
Moreover, we can find that $g$ is nonincreasing with respect to $x$. From the Proposition 9.1.2 in [4], we have
$\mathbb{E}[g(\min_{i\in\{1,2,...n\}}\{X_{i}+Y_{i}\})]\leq\mathbb{E}[g(\min_{i\in\{1,2,...n\}}\{S_{i}+T_{i}\})].$ (12)
We define $W=\frac{1}{\min_{i\in\{1,2,...n\}}\{S_{i}+T_{i}\}}$, then $W\in[\frac{\phi}{2},\infty)$. Since $S_{i},T_{i}$ are independent uniformly distributed random variables on $[0,\frac{1}{\phi}]$, we have
$F_{W}(w)=Prob[W\leq w]=\prod^{n}_{i=1}Prob[S_{i}+T_{i}\geq\frac{1}{w}]=(1-\frac{\phi^{2}}{2w^{2}})^{n}.$
For $i=0,1,...,\lambda$, we define $Q_{i}$ as the random event that $W\in[\frac{\phi}{2}\cdot n^{i},\frac{\phi}{2}\cdot n^{i+1})$. Moreover, we let $Q_{\lambda+1}$ denote the random event that $W\in[\frac{\phi}{2}\cdot n^{\lambda+1},\infty)$. Then, we can upper bound the expectation in (12) by:
$\mathbb{E}[g(\min_{i\in\{1,2,...n\}}\{X_{i}+Y_{i}\})]$ $\leq\mathbb{E}[\min\{\alpha\cdot Wln(\alpha\cdot W),n^{\lambda}\}]$
$\leq\sum^{\lambda}_{i=0}Prob[Q_{i}]\cdot\mathbb{E}[\alpha\cdot Wln(\alpha\cdot W)]+Prob[Q_{\lambda+1}]\cdot n^{\lambda}$
$\leq\sum^{\lambda}_{i=0}[F_{W}(\frac{\phi}{2}\cdot n^{i+1})-F_{W}(\frac{\phi}{2}\cdot n^{i})]\cdot\frac{\alpha\phi}{2}\cdot n^{i+1}\cdot ln(\frac{\alpha\phi}{2}\cdot n^{i+1})$
$\,\quad+[1-F_{W}(\frac{\phi}{2}\cdot n^{\lambda+1})]\cdot n^{\lambda}$
$\leq\frac{\alpha\phi}{2}\cdot ln(\frac{\alpha\phi n}{2})\cdot\sum^{\lambda}_{i=0}[F_{W}(\frac{\phi}{2}\cdot n^{i+1})-F_{W}(\frac{\phi}{2}\cdot n^{i})]\cdot n^{i+1}\cdot(i+1)$
$\,\quad+[1-F_{W}(\frac{\phi}{2}n^{\lambda+1})]\cdot n^{\lambda}$
$=\frac{\alpha\phi}{2}\cdot ln(\frac{\alpha\phi n}{2})\cdot\sum^{\lambda}_{i=0}[(1-\frac{2}{n^{2i+2}})^{n}-(1-\frac{2}{n^{2i}})^{n}]\cdot n^{i+1}\cdot(i+1)$
$\,\quad+[1-(1-\frac{2}{n^{2\lambda+2}})^{n}]\cdot n^{\lambda}$
$\leq\alpha\phi n\cdot ln(\alpha\phi n)\cdot(\lambda+1)^{2}.$
Therefore, we complete the proof of the lemma.
Lemma 3
Let $x\in(0,1)$, then we have:
$-\frac{1}{ln(1-x)}\leq\frac{1}{x}$
Proof
Since $x\in(0,1)$, we have $-ln(1-x)\geq 0$. Moreover, it is easy to find out that $ln(1-x)=ln(\frac{1}{1-x})$ and $x=1-\frac{1}{\frac{1}{1-x}}$. By defining $t=\frac{1}{1-x}$, we only need to prove
$lnt\geq 1-\frac{1}{t}$
for any $t>1$. We define $f(t)=lnt+\frac{1}{t}-1$, then $f^{\prime}(t)=\frac{t-1}{t^{2}}\geq 0$. Therefore, $f(t)\geq f(1)=0$, which completes the proof.
Now we begin to prove that $\epsilon$-ABRD always find the $\alpha$-APNE in a polynomial time. It is easy to see that we only need to prove the expected number of iterations is polynomial in $\frac{1}{\epsilon},\phi,n,m$.
@@
Given an instance $(G,\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$ of the game and $c_{e}(W_{e})=a_{e}W_{e}+b_{e}$, where $\{a_{e},b_{e}\}_{e\in E}$ are $\phi$-smooth random variables. Then, the expected number of iterations of $\epsilon$-ABRD is at most $O(\frac{1}{\epsilon}W^{3}\phi n^{3}m^{3}ln(W\phi nm))$.
@@
$\Phi(P)\leq\sum_{e\in E}c_{e}(W)log_{2}(1+W_{e})\leq m\max_{e\in E}\{a_{e}W+b_{e}\}log_{2}(1+W).$ (13)
To simplify our discussion, we define $c_{max}=\max_{e\in E}\{a_{e}W+b_{e}\}$. For the lower bound of the approximate potential function $\Phi$, we first propose a fact as following:
$log_{2}(1+x)+log_{2}(1+y)\geq log_{2}(1+x+y)$
for any $x,y>0$. And since $\min_{i\in\mathcal{N}}w_{i}=1$ and each player uses at least one edge, then $\Phi$ can be lower bounded by
$\Phi(P)\geq\sum_{e\in E}\min_{e\in E}\{a_{e}+b_{e}\}log_{2}(1+W_{e})\geq\min_{e\in E}\{a_{e}+b_{e}\}log_{2}(1+W).$ (14)
To simplify our discussion, we define $c_{min}=\min_{e\in E}\{a_{e}+b_{e}\}$.
@@
$\geq\frac{\epsilon c_{min}}{(1+\epsilon)\cdot W\cdot log_{2}(1+W)\cdot m\cdot c_{max}}\Phi(P).$ (15)
So we have
$\Phi(P^{\prime})<(1-\epsilon^{\prime})\Phi(P),\quad\text{where}\;\epsilon^{\prime}=\frac{\epsilon c_{min}}{(1+\epsilon)Wlog_{2}(1+W)mc_{max}}.$ (16)
@@
$c_{min}log_{2}(1+W)\leq\Phi(P^{T})<(1-\epsilon^{\prime})^{T}\Phi(P^{0})\leq(1-\epsilon^{\prime})^{T}mc_{max}log_{2}(1+W).$
@@
Since $10<\epsilon^{\prime}<$ 1, we can derive the following upper bound on the number of steps for our $\epsilon$-ABRD from Lemma 3 and Inequality (17):
@@
$=(1+\frac{1}{\epsilon})Wlog_{2}(1+W)\frac{mc_{max}}{c_{min}}ln\frac{mc_{max}}{c_{min}}.$ (18)
@@
$T\leq\min\{(1+\frac{1}{\epsilon})Wlog_{2}(1+W)\frac{mc_{max}}{c_{min}}ln\frac{mc_{max}}{c_{min}},2^{nm}\}.$ (19)
@@
$\mathbb{E}[T]$ $\leq(1+\frac{1}{\epsilon})Wlog_{2}(1+W)\mathbb{E}[\min\{\frac{m(1+W)}{c_{min}}ln\frac{m(1+W)}{c_{min}},n^{nmlog_{n}2}\}]$
$\leq(1+\frac{1}{\epsilon})Wlog_{2}(1+W)\cdot m(1+W)\phi\cdot n\cdot ln(m(1+W)\phi\cdot n)(nmlog_{n}2+1)^{2}$
$\leq 4\cdot(1+\frac{1}{\epsilon})W^{3}\cdot\phi\cdot n^{3}\cdot m^{3}\cdot ln(W\cdot\phi\cdot n\cdot m)$
$=O(\frac{1}{\epsilon}W^{3}\cdot\phi\cdot n^{3}\cdot m^{3}\cdot ln(W\cdot\phi\cdot n\cdot m)),$

### 17 Dec 2023 v2

@@ A $(1+\epsilon)\alpha$-APNE of game $\Gamma$
    While $P$ is not an $(1+\epsilon)\alpha$-APNE do
@@
Specifically, the first and the second line of the Algorithm 1 is finding an $(1+\epsilon)\alpha$-improving deviation for player $i$. For a given instance of game $\Gamma$, we may not be able to search exhaustively in the strategy space $\mathcal{P}_{i}$ since it will cost exponential time. Instead, we can make a shortest-path search for player $i$, which can find the best response for player $i$ efficiently and cost polynomial time. In this way, we only need to verify whether the best response is an $(1+\epsilon)\alpha$-improving deviation, which can also be computed in polynomial time. Therefore, in order to show the Algorithm 1 runs in polynomial time, we only need to prove the number of iterations is polynomial in the given parameters. Before we begin to prove, we first make an assumption of the parameter and propose a lemma as preparation.
@@
Let $X$ be a random variable, we say $X$ is a $\phi$-smooth random variable if and only if $X$ is defined on $[0,1]$ and the probability density function of $X$ is up to $\phi$($\phi\geq 1$).
In our model, it is natural to assume $\{a_{e},b_{e}\}_{e\in E}$ are $\phi$-smooth random variables. For $\phi$-smooth random variable, we can derive the lemma below.
@@
Let $X_{1},X_{2},...,X_{n};Y_{1},Y_{2},...,Y_{n}(n\geq 2)$ be $\phi$-smooth independent random variables on [0,1]. Let $g(x)=\min\{\frac{\alpha}{x}ln\frac{\alpha}{x},n^{\lambda}\}$. Then for any $\lambda\in N$ and $\alpha\geq 1$, we have
$\mathbb{E}[g(\min_{i\in[n]}\{X_{i}+Y_{i}\})]\leq\alpha\phi(n^{2}+1)\cdot ln(\alpha\phi n)$ (7)
@@
For two random variables $S,T$, we define $S\preceq T$ if and only if $F_{T}(x)\leq F_{S}(x)$ for any $x\in\mathbb{R}$, where $F_{S}(x),F_{T}(x)$ are the cumulative distribution function of $S,T$ respectively. To prove the lemma, we define two sets of independent uniformly distributed random variables $S_{1},S_{2},...,S_{n};T_{1},T_{2},...,T_{n}$ on $[0,\frac{1}{\phi}]$. In addition, we denote $[n]=\{1,2,...n\}$ to simplify our discussion.
@@
For any $i\in[n]$, we have
$min_{i\in[n]}S_{i}\preceq min_{i\in[n]}X_{i},min_{i\in[n]}T_{i}\preceq min_{i\in[n]}Y_{i}.$
@@
Since $S_{i}$ is defined on $[0,\frac{1}{\phi}]$, we have $F_{S_{i}}(t)=1$ for any $t\geq\frac{1}{\phi}$, then we have $F_{X_{i}}(t)\leq F_{S_{i}}(t)$ definitely. When $0\leq t\leq\frac{1}{\phi}$, since $X_{i}$ is a $\phi$-smooth random variable for any $i\in[n]$, its probability density function is upper-bounded by $\phi$. Then we have
@@
Moreover, we can find that $g(x)=\min\{\frac{\alpha}{x}ln\frac{\alpha}{x},n^{\lambda}\}$ is nonincreasing with respect to $x$. From the Proposition 9.1.2 in [4], we have
$\mathbb{E}[g(\min_{i\in[n]}\{X_{i}+Y_{i}\})]\leq\mathbb{E}[g(\min_{i\in[n]}\{S_{i}+T_{i}\})].$ (8)
We define $W=\frac{1}{\min_{i\in[n]}\{S_{i}+T_{i}\}}$, then $W\in[\frac{\phi}{2},\infty)$. Since $S_{i},T_{i}$ are independent uniformly distributed random variables on $[0,\frac{1}{\phi}]$, we have
$F_{W}(w)=\prod^{n}_{i=1}Prob[S_{i}+T_{i}\geq\frac{1}{w}]=\left\{\begin{aligned} &(1-\frac{\phi^{2}}{2w^{2}})^{n},\quad\quad\quad\ \ w\in[\phi,\infty)\\ &(2-\frac{2\phi}{w}+\frac{\phi^{2}}{2w^{2}})^{n},\quad w\in[\frac{\phi}{2},\phi)\end{aligned}.\right.$ (9)
The last equation is due to Lemma 3 where we let $A\leftarrow\frac{1}{\phi},t\leftarrow\frac{1}{w}$.
For $i=1,2...,\lambda,\lambda+1$, we define $Q_{i}$ as the random event that $W\in[\phi n^{i-1},\phi n^{i})$. Moreover, we let $Q_{0}$ and $Q_{\lambda+2}$ denote the random event that $W\in[\frac{\phi}{2},\phi)$ and $W\in[\phi n^{\lambda+1},\infty)$ respectively. Then, we can derive the following upper bound on the expectation in Inequality (8) by:
$\mathbb{E}[g(\min_{i\in[n]}\{X_{i}+Y_{i}\})]$ $\leq\mathbb{E}[\min\{\alpha\cdot Wln(\alpha\cdot W),n^{\lambda}\}]$
$\leq\sum^{\lambda+1}_{i=0}Prob[Q_{i}]\cdot\mathbb{E}[\alpha\cdot Wln(\alpha\cdot W)|Q_{i}]+Prob[Q_{\lambda+2}]\cdot n^{\lambda}$
$\leq\sum^{\lambda+1}_{i=1}[F_{W}(\phi n^{i})-F_{W}(\phi n^{i-1})]\cdot\alpha\phi n^{i}\cdot ln(\alpha\phi n^{i})$
$\,\quad+[F_{W}(\phi)-F_{W}(\frac{\phi}{2})]\alpha\phi ln(\alpha\phi)+[1-F_{W}(\phi n^{\lambda+1})]n^{\lambda}$
$\leq\alpha\phi\cdot ln(\alpha\phi n)\cdot\sum^{\lambda+1}_{i=1}[F_{W}(\phi n^{i})-F_{W}(\phi n^{i-1})]\cdot n^{i}\cdot i$
$\,\quad+[F_{W}(\phi)-F_{W}(\frac{\phi}{2})]\alpha\phi ln(\alpha\phi)+[1-F_{W}(\phi n^{\lambda+1})]n^{\lambda}$
$=\alpha\phi\cdot ln(\alpha\phi n)\cdot\sum^{\lambda+1}_{i=1}[(1-\frac{1}{2n^{2i}})^{n}-(1-\frac{1}{2n^{2i-2}})^{n}]\cdot in^{i}$
$\,\quad+(\frac{1}{2})^{n}\cdot\alpha\phi\cdot ln(\alpha\phi)+[1-(1-\frac{1}{2n^{2\lambda+2}})^{n}]\cdot n^{\lambda}$
$\leq\alpha\phi\cdot ln(\alpha\phi n)\cdot\sum^{\lambda+1}_{i=1}[\frac{1}{1+\frac{n}{2n^{2i}}}-(1-\frac{n}{2n^{2i-2}})]\cdot in^{i}$
$\,\quad+(\frac{1}{2})^{n}\cdot\alpha\phi\cdot ln(\alpha\phi)+[1-(1-\frac{n}{2n^{2\lambda+2}})]\cdot n^{\lambda}$
$\leq\alpha\phi\cdot ln(\alpha\phi n)\cdot\sum^{\lambda+1}_{i=1}\frac{i}{2n^{i-3}}+(\frac{1}{2})^{n}\cdot\alpha\phi\cdot ln(\alpha\phi)+\frac{1}{2n^{\lambda+1}}$
$\leq\alpha\phi\cdot ln(\alpha\phi n)\cdot(n^{2}+1).$
The second inequality is due to the law of total expectation. And the third to last inequality is due to Lemma 2. Therefore, we complete the proof of the lemma.
It is worth mentioning that there is no $\lambda$ on the right side of Inequality 7. With the formula for changing base, the second term of the function $g$ has almost no effect on the conclusion.
Now we begin to show that $\epsilon$-ABRD always find the $(1+\epsilon)\alpha$-APNE in a polynomial time by proving the expected number of iterations is polynomial in the given parameters.
@@
Given an instance $(G,\mathcal{N},\mathcal{P},\mathcal{C},\mathcal{W})$ of the game and $c_{e}(w_{e})=a_{e}w_{e}+b_{e}$, where $\{a_{e},b_{e}\}_{e\in E}$ are $\phi$-smooth random variables. Then, the expected number of iterations of $\epsilon$-ABRD is at most $O(\frac{1}{\epsilon}W^{3}\phi n^{2}m\cdot ln(W\phi nm))$.
@@
$\Phi(P)\leq\sum_{e\in E}c_{e}(W)\log_{2}(1+w_{e})\leq m\max_{e\in E}\{a_{e}W+b_{e}\}\log_{2}(1+W).$ (10)
To simplify our discussion, we let $c_{max}=\max_{e\in E}\{a_{e}W+b_{e}\}$. Since $\min_{i\in\mathcal{N}}w_{i}=1$ and each player uses at least one edge, then $\Phi$ can be lower bounded by
$\Phi(P)\geq\sum_{e\in E}\min_{e\in E}\{a_{e}+b_{e}\}\log_{2}(1+w_{e})\geq\min_{e\in E}\{a_{e}+b_{e}\}\log_{2}(1+W).$ (11)
The second inequality is due to $\log_{2}(1+x)+\log_{2}(1+y)\geq\log_{2}(1+x+y)$ for any $x,y>0$. To simplify our discussion, we let $c_{min}=\min_{e\in E}\{a_{e}+b_{e}\}$.
@@
$\geq\frac{\epsilon c_{min}}{(1+\epsilon)\cdot W\cdot\log_{2}(1+W)\cdot m\cdot c_{max}}\Phi(P).$ (12)
Now we have
$\Phi(P^{\prime})<(1-\epsilon^{\prime})\Phi(P),\quad\text{where}\;\epsilon^{\prime}=\frac{\epsilon c_{min}}{(1+\epsilon)W\log_{2}(1+W)mc_{max}}.$ (13)
@@
$c_{min}\log_{2}(1+W)\leq\Phi(P^{T})<(1-\epsilon^{\prime})^{T}\Phi(P^{0})\leq(1-\epsilon^{\prime})^{T}mc_{max}\log_{2}(1+W).$
@@
Since $10<\epsilon^{\prime}<$ 1, we can derive the following upper bound on the number of steps for our $\epsilon$-ABRD from Inequality (14):
@@
$=(1+\frac{1}{\epsilon})W\log_{2}(1+W)\frac{mc_{max}}{c_{min}}ln\frac{mc_{max}}{c_{min}}.$ (15)
The second inequality is due to $\epsilon^{\prime}\geq ln(\frac{1}{1-\epsilon^{\prime}})$ when $10<\epsilon^{\prime}<$ 1.
@@
$T\leq\min\{(1+\frac{1}{\epsilon})W\log_{2}(1+W)\frac{mc_{max}}{c_{min}}ln\frac{mc_{max}}{c_{min}},2^{nm}\}.$ (16)
@@
$\mathbb{E}[T]$ $\leq(1+\frac{1}{\epsilon})W\log_{2}(1+W)\mathbb{E}[\min\{\frac{m(1+W)}{c_{min}}ln\frac{m(1+W)}{c_{min}},m^{nm}\}]$
$\leq(1+\frac{1}{\epsilon})W\log_{2}(1+W)\cdot m(1+W)\phi(n^{2}+1)\cdot ln(m(1+W)\phi\cdot n)$
$\leq 4\cdot(1+\frac{1}{\epsilon})W^{3}\cdot\phi\cdot n^{2}\cdot m\cdot ln(W\cdot\phi\cdot n\cdot m)$
$=O(\frac{1}{\epsilon}W^{3}\cdot\phi\cdot n^{2}\cdot m\cdot ln(W\cdot\phi\cdot n\cdot m)),$
@@
5 Conclusion
Our paper give the first result on weighted Shapley network design game with affine linear cost function $c_{e}(w_{e})=a_{e}w_{e}+b_{e}$. The first major result is the existence of $O(\log_{2}(W))$-APNE, which is shown by constructing an $O(\log_{2}(W))$-approximate potential function $\Phi(P)=\sum_{e\in E}c_{e}(w_{e})\log_{2}(1+w_{e})$. For the weighted Shapley network design game with constant cost function $c_{e}$, Chen and Roughgarden [7] proved the existence of $\log_{2}[e(1+w_{max})]$-APNE. It can be seen from the comparison that the value of $\alpha$ becomes a little larger but is still within the acceptable range.
On the other hand, computing $(1+\epsilon)\alpha$-APNE in polynomial time through $\epsilon$-best response dynamics is another major result. Since we can adopt the shortest-path search to ensure every step of iteration runs in polynomial time, we only need to prove the expected number of iterations is polynomial in parameters. Specifically, We assume $\{a_{e},b_{e}\}_{e\in E}$ are $\phi$-smooth random variables and prove the expected number of iterations is at most $O(\frac{1}{\epsilon}W^{3}\phi n^{2}m\cdot ln(W\phi nm))$, which completes the analysis on the time complexity of $\epsilon$-best response dynamics. For the unweighted Shapley network design game with $\phi$-smooth random variables $c_{e}$, Giannakopoulos [12] proved the expected number of iterations of $\epsilon$-best response dynamics is $O(\frac{1}{\epsilon}\phi\log(\phi)n\log^{3}(n)m^{5})$. Then we can find out that the number of $m$ in the expected number of iterations is greatly reduced in our model.
In the future, there are plenty of issues with the model that are worth studying, such as the price of anarchy and the price of stability of the game. Furthermore, we can further generalize the cost function into a polynomial function or even a general continuous function and study the $\alpha$-APNE of the game.
Acknowledgements
This work was partially supported by the National Natural Science Foundation of China (No. 12161141006), the Natural Science Foundation of Tianjin (No. 20JCJQJC00090).
Appendix
Lemma 2
Let $x\in(0,1),n\in N$, then we have:
  1. 1.
    $(1-x)^{n}\geq 1-nx$;
  2. 2.
    $(1-x)^{n}\leq\frac{1}{1+nx}$.
Proof
We use mathematical induction to prove part 1. When $n=1$, the left and right sides of the inequality are equal. If the inequality holds when $n=k$, then for $n=k+1$, we have
$(1-x)^{k+1}$ $=(1-x)^{k}(1-x)$
$>(1-kx)(1-x)$
$=1-(k+1)x+kx^{2}$
$>1-(k+1)x.$
Therefore, by mathematical induction we prove part 1.
For part 2, from binomial expansion we have $(1+x)^{n}\geq 1+nx$. Then we can derive the following:
$1>(1-x^{2})^{n}=(1+x)^{n}(1-x)^{n}\geq(1+nx)(1-x)^{n}$
Therefore, we prove $(1-x)^{n}\leq\frac{1}{1+nx}$ and complete the proof of lemma.
Lemma 3
Let $X,Y$ be two independent uniformly distributed random variables on $[0,A]$, where $A>0$ is a constant. For a random variable $X$, $F_{X}(t)$ denotes the cumulative function of $X$. Then we have:
$F_{X+Y}(t)=\left\{\begin{aligned} &\frac{t^{2}}{2A^{2}},\qquad\qquad\quad t\in[0,A)\\ &\frac{2t}{A}-\frac{t^{2}}{2A^{2}}-1,\quad t\in[A,2A]\end{aligned}.\right.$
Proof
We assume $Z=X+Y$, then $Z$ is defined on $[0,2A]$. For a random variable $X$, let $f_{X}(t)$ be the probability density function of $X$. Then we have $f_{X}(t)=f_{Y}(t)=\frac{1}{A}$ for any $t\in[0,A]$. Since $X,Y$ are independent random variables, we have
[…]
When $t\in[0,A)$, from Equation 17 we have
[…]
Similarly, we can derive the following from Equation 17 when $t\in[A,2A]$:
[…]
Furthermore, we have
$F_{Z}(t)=\int^{t}_{0}f_{Z}(x)dx=\left\{\begin{aligned} &\frac{t^{2}}{2A^{2}},\qquad\qquad\quad t\in[0,A)\\ &\frac{2t}{A}-\frac{t^{2}}{2A^{2}}-1,\quad t\in[A,2A]\end{aligned}.\right.$